# Double Privacy Guard: Robust Traceable Adversarial Watermarking against Face Recognition

### v1

@@ IV-A Information-Guided Adversarial Attack
To enhance the generalization capability of the DPG encoder and decoder and avoid overfitting, we generate a random watermark for each batch during training. Users can customize $w_{id}$ for one or more images during the inference process. To limit the impact of watermark encoding on the carrier, we use MSE loss $\mathcal{L}_{MSE}$ to define the image invisibility loss $\mathcal{L}_{inv}^{\phi_{en}}$:
@@
To make the traceable watermark adversarial against FR models, we introduce adversarial constraints to optimize the encoder. FR model attacks can be categorized into two types: untargeted (dodging) and targeted (impersonation) attacks. Unlike dodging attack lowering similarity confidence for same-identity pairs, impersonation attack aims to enhance feature similarity between target and source images of different identities [10]. This method seamlessly incorporate the watermark model into end-to-end optimization and disrupt the feature reconstruction of FR models. Therefore, we focus on impersonation attacks as in [10, 4, 1].
To achieve the mapping of recognizable features from the watermarked image to the target domain, we first obtain recognizable feature embeddings $E_{\ast}$ from the FR model for both the watermarked image and the target image, as follows:
@@
IV-B Muti-Task Compatible Meta-Optimization
The DPG encoder is associated with two distinct sub-tasks: 1) encoding the watermark recoverably and invisibly; 2) adversarial encoding the watermark to deceive FR models. Simultaneously, the decoder needs to accurately extract the watermark. DPG are optimized end-to-end to search for the global optimal solution. However, there are optimization conflicts between different tasks. Enhancing adversariality only under a single FR model or after simply ensemble FR models hinders watermark invisibility optimization, introducing more obvious perceptual artifacts to the watermarked image. Additionally, this adversarial enhancement is not effectively transferred to black-box models. Therefore, to enhance the transferability and mitigate this optimization conflict, we adopt a collaborative meta-optimization strategy compatible with multi-task. This strategy builds upon Model-Agnostic Meta-Learning (MAML) method [22] and further corrects the training errors of different tasks, seamlessly integrating into end-to-end training to aid overall optimization. The schematic diagram of model parameter updates in the overall optimization is shown in Figure 3.
We select $P$ meta-training models and $Q$ meta-testing models, initially calculating the adversarial loss under the meta-training models with the initial encoder parameters $\phi_{en}$:
$\mathcal{L}_{adv}^{tra}=\sum_{p=1}^{P}\mathcal{L}_{adv}(\mathcal{F}_{p}(\hat{x}_{s}),\mathcal{F}_{p}(x_{t})\mid\phi_{en}).$ (8)
@@ IV-B Muti-Task Compatible Meta-Optimization
IV-C Regularization-Enhanced Information Decoding
Online networks process uploaded images to conform to file storage and transmission standards. This processing significantly changes the watermark distribution and increases the difficulty of watermark extraction. Therefore, we design a joint noise pool in front of the decoder, which includes differentiable JPEG compression ($QF\!\!=\!\!50$) and Gaussian noise ($Var\!\!=\!\!0.003$), enhancing the decoder robustness. We initially calculate the information loss of the watermarked image obtained under the original encoder parameters using BCE loss $\mathcal{L}_{BCE}$, as follows:
@@ IV-C Regularization-Enhanced Information Decoding
The encoder and decoder interact synergistically during co-optimization, serving as mutual priors, as illustrated in Figure 3. To enhance cross-model gradient compatibility, we introduce an information regularization module. This module calculates information loss using the latest temporary update of the watermarked image, allowing it to further refine the optimization direction of the decoder based on the encoder:
@@
The outer loop optimization of the model updates the encoder and decoder with adversarial loss, invisibility loss, and watermark information loss as joint optimization objectives. The total loss of DPG is calculated as follows:
@@
where $\lambda_{adv}$, $\lambda_{en}$ and $\lambda_{wm}$ are positive hyperparameters, and we empirically set them to 100, 0.05 and 0.05, respectively.
V Experiments
V-A Experimental Settings
V-A1 Datasets
We evaluate DPG on two commonly used large face image datasets: LFW [23] and CelebA [24]. The watermark message is 50 bits, sufficient to meet the requirements for representing identity in the real world. To meet the input size requirements of FR models, all images are initially processed through MTCNN [25] to extract facial regions and sampled to a size of $112\times 112$.
V-A2 Implementation details
The entire training process of DPG spans 2500 epochs with a batch size of 32. We empirically adjust the Adam optimizer with an initial learning rate of 0.00005 for stable training. We select various offline and online commercial victim FR models to evaluate the black-box transferability of DPG,including: 1) Offline models: IR152, IRSE50, MobileFace [20], FaceNet [26], ArcFace [20] based on IResNet100 [27] and CosFace [28] based on IResNet100. 2) Online models: Face++ [29] and Aliyun [30]. We alternately selected three models from them as white-box models for integrated training and sequentially obtained black-box test results on the remaining models as in [10, 9, 31], with the parameters set as $P=2$ and $Q=1$.
V-A3 Comparisons
DPG is the first watermarking method designed for FR models with both adversariality and traceability, and there is no similar work before. Thus, we first compare its adversariality with adversarial example methods: 1) Patch-based methods: Adv-Makeup [10], AdvHat [9]; 2) Perturbation-based methods: PGD [11], FGSM [12]. Simultaneously, we straightforwardly combine adversarial example methods with watermarking methods Hidden [7] and FakeTagger [14]. We embed watermarks directly into adversarial examples generated by PGD and FGSM, comparing adversariality and traceability to illustrate the advantages of DPG.
  Datasets Type Method IR152 IRSE50 ArcFace MobileFace FaceNet CosFace
@@ V-A3 Comparisons
LFW Patch-based Adv-Makeup 0.026 N/A 0.236 N/A 0.026 N/A 0.289 N/A 0.026 N/A 0.00 N/A
@@
Watermark-based DPG (ours) 0.326 0.992 0.810 0.997 0.381 0.996 0.716 0.997 0.643 0.994 0.355 0.996
@@
Watermark-based DPG (ours) 0.433 0.998 0.635 0.998 0.375 0.995 0.679 0.997 0.318 0.995 0.393 0.998
@@
Watermark-based DPG (ours) 65.491 0.998 43.945 0.998
@@
V-A4 Metrics
We use SSIM [32] to evaluate the quality of images. Accuracy $Acc$ is used to evaluate the watermark recovery accuracy of the watermarked image. we adopt Attack Success Rate (ASR) [33] for impersonation attack to evaluate adversariality, which is calculated through:
@@ V-A4 Metrics
where $i$ represents the image index, $N$ represents the total number of test images, and $\tau$ represents the similarity threshold. The value of $\tau$ will be set as the threshold at 0.01 FAR (False Acceptance Rate) for each victim FR model as most face recognition works do, i.e., IR152 (0.167), IRSE50 (0.241), MobileFace (0.302), Facenet (0.409), ArcFace (0.280) and CosFace (0.280) [10].
V-B Comparison with Baselines
To meet the challenges of the practical scenarios, where users may lack knowledge of the underlying structure and the specific training schemes of the FR models used by third-party databases, we focus on evaluating the defense method’s performance against black-box models. To ensure that the generated adversarial images do not affecting normal usage, we controlled the average SSIM values of all adversarial images from different methods to be 0.9 $\pm$ 0.03 in our experiments for fair comparison. As shown in Table II, DPG exhibits excellent adversariality and traceability simultaneously. It achieves optimal ASR on almost all evaluated black-box FR models, surpassing adversarial example methods that only possess adversariality. Moreover, DPG outperforms combination-based methods in both adversariality and traceability, especially on FaceNet, ArcFace, and CosFace, where watermark embedding significantly reduces the adversariality of the generated examples. Most importantly, DPG achieves two significant performances by training only one set of watermarking models, providing a more unified management approach for users or authorizers over models and images.
To assess the practical viability of the attack methods, we leverage the face comparison feature of online commercial FR models. We compared the generated adversarial images from different methods with the target facial images, and the platform provided confidence score indicating whether the two images belong to the same identity. We reported the Average Confidence Scores (ACS) for 100 images from the LFW dataset. As shown in Table III, DPG achieved the highest ACS while maintaining excellent traceability accuracy. This confirms the feasibility of DPG for practical deployment.
  Processing Parameters Type Method IR152 IRSE50 ArcFace MobileFace FaceNet CosFace
@@ V-B Comparison with Baselines
GN 0.003 Patch-based Adv-Makeup 0.026 N/A 0.263 N/A 0.026 N/A 0.289 N/A 0.026 N/A 0.000 N/A
@@
Combination-based PGD$+$Hidden 0.211 0.629 0.424 0.629 0.055 0.632 0.365 0.629 0.033 0.632 0.016 0.632
FGSM$+$Hidden 0.113 0.633 0.145 0.633 0.000 0.636 0.312 0.633 0.011 0.636 0.000 0.636
PGD$+$FakeTagger 0.220 0.971 0.410 0.971 0.023 0.961 0.341 0.971 0.017 0.961 0.022 0.961
FGSM$+$FakeTagger 0.103 0.970 0.312 0.970 0.013 0.963 0.300 0.970 0.011 0.963 0.024 0.963
Watermark-based DPG (ours) 0.265 0.975 0.725 0.987 0.308 0.979 0.695 0.987 0.563 0.975 0.311 0.985
JPEG 30 Patch-based Adv-Makeup 0.000 N/A 0.052 N/A 0.000 N/A 0.131 N/A 0.000 N/A 0.000 N/A
@@
Combination-based PGD$+$Hidden 0.261 0.654 0.513 0.654 0.032 0.662 0.412 0.654 0.030 0.662 0.045 0.662
FGSM$+$Hidden 0.119 0.686 0.280 0.686 0.000 0.655 0.295 0.686 0.035 0.655 0.016 0.655
PGD$+$FakeTagger 0.260 0.886 0.463 0.886 0.040 0.883 0.440 0.898 0.021 0.883 0.029 0.883
FGSM$+$FakeTagger 0.119 0.887 0.310 0.887 0.016 0.883 0.360 0.887 0.022 0.883 0.011 0.883
Watermark-based DPG (ours) 0.272 0.967 0.655 0.958 0.286 0.983 0.736 0.967 0.587 0.983 0.398 0.980
Resize $1/2$ Patch-based Adv-Makeup 0.026 N/A 0.236 N/A 0.026 N/A 0.289 N/A 0.026 N/A 0.000 N/A
@@
Combination-based PGD$+$Hidden 0.241 0.691 0.411 0.691 0.023 0.641 0.444 0.691 0.022 0.641 0.023 0.641
FGSM$+$Hidden 0.109 0.714 0.312 0.714 0.000 0.721 0.333 0.714 0.011 0. 721 0.000 0.721
PGD$+$FakeTagger 0.203 0.986 0.336 0.986 0.025 0.988 0.425 0.986 0.011 0.988 0.016 0.988
FGSM$+$FakeTagger 0.137 0.987 0.246 0.987 0.000 0.986 0.361 0.987 0.011 0.986 0.000 0.986
Watermark-based DPG (ours) 0.255 0.990 0.584 0.990 0.211 0.989 0.688 0.990 0.408 0.993 0.253 0.988
@@
V-C Robustness Analysis
Online platforms often process images to conform to uniform transmission standards. Therefore, we evaluate the adversarial robustness and traceability robustness of the methods after applying Gaussian noise (GN), JPEG compression (JPEG), or resizing (Resize), as shown in Table IV. DPG maintains optimal performance even after various image processing, surpassing dedicated adversarial attack methods by up to $58\%$ in adversarial robustness and outperforming combination-based methods in traceability robustness. This robustness advantage can be attributed to two factors. Firstly, unlike adversarial attacks that directly overlay gradients, DPG embeds watermarks in the deep feature space of the carrier. Additionally, DPG synergistically optimizes the encoder and decoder, introducing adversarial attack to increase the output dimensions of the model.  [21] has confirmed that multitask learning, especially the joint optimization of tasks with low correlation, significantly enhances model robustness.
V-D Visualization and Analysis
We present the visual results of different methods using samples covering various genders, ages, and degrees of facial occlusion, as shown in Figure 4. Patch-based methods rely on significant semantic modifications to images. Adv-Makeup has limitations for images with occluded eyes. Perturbation-based methods produce adversarial examples with regular noise patterns, easy to be discovered and defended by adversaries. DPG-generated watermarked images exhibit only differences in overall hue, without noticeable artifacts. Additionally, we use Grad-CAM [34] visualization to illustrate the gradient responses of FGSM and DPG under black-box and white-box models. As shown in Figure 5, FGSM and DPG both modify gradient responses in the white-box model. While FGSM overfit to modifications in the region of interest of the white-box model, including the background, resulting in poor transferability to black-box models. DPG encodes more universal adversarial features of the carrier, focusing on crucial facial regions, and effectively transfers these adversariality to the black-box model.
  Method IR152 IRSE50 ArcFace MobileFace FaceNet CosFace

### v2

@@ 4.1 Information-Guided Adversarial Attack
To enhance the generalization capability of the DIP-Watermark encoder and decoder and avoid overfitting, we generate a random watermark for each batch during training process. In practical use, the platform registers an identity-specific watermark for each user, and the watermark information can be customized by the platform according to the users. To limit the impact of watermark encoding on the carrier, we use MSE loss $\mathcal{L}_{MSE}$ to define the image invisibility loss $\mathcal{L}_{inv}^{\phi_{en}}$:
@@
To make the traceable watermark adversarial against FR models, we introduce adversarial constraints to optimize the encoder. FR model attacks can be categorized into two types: untargeted (dodging) and targeted (impersonation) attacks. dodging attacks aim to reduce the similarity of images of the same identity to attack the FR model, while impersonation attacks aim to enhance feature similarity between target and source images of different identities [16], thus disrupting the decision of the FR model. Dodging attacks hinder the stable training of the watermark model, easily falling into a local optimum of watermarked image quality optimization in the early stages of training. Impersonation attacks can seamlessly integrate into the end-to-end optimization process of the watermark model, unifying the multi-task optimization direction of the watermark, and disrupting the feature reconstruction of the FR model. Therefore, we focus on impersonation attacks as in [16, 8, 1].
We map the identifiable features of the source image to the target identity domain by optimizing the embedding of watermark information. we first obtain recognizable feature embeddings $E_{\ast}$ from the FR model for both the watermarked image and the target image, as follows:
@@
4.2 Muti-Task Compatible Meta-Optimization
The DIP-Watermark encoder is associated with two distinct sub-tasks: 1) encoding the watermark recoverably and invisibly; 2) adversarial encoding the watermark to deceive FR models. Simultaneously, the decoder needs to accurately extract the watermark. DIP-Watermark are optimized end-to-end to search for the global optimal solution. However, during the watermark model training process, there are optimization conflicts between different tasks. Enhancing the watermark adversariality on a single FR model or on a simple ensemble of FR models would lead the encoder to blindly enhance the watermark embedding strength, thus hindering the optimization of watermark invisibility, introducing more noticeable perceptual artifacts into the watermarked image. More importantly, such adversariality enhancement cannot be effectively transferred to black-box models. Therefore, to enhance the transferability and mitigate this optimization conflict, we introduce a collaborative meta-optimization strategy compatible with multi-task. This strategy builds upon Model-Agnostic Meta-Learning (MAML) method [28]. MAML encodes more representative shared adversarial features from the carrier by combining different known models, thus quickly adapting to unknown target models. We seamlessly integrate adversarial transferability optimization into other subtasks of the encoder to aid the overall optimization of the watermark model.The schematic diagram of model parameter updates in the overall optimization is shown in Fig. 3.
In the meta-optimization strategy, the known target victim models are divided into meta-training models and meta-testing models. Temporary model parameters are calculated using the loss obtained on the meta-training models, and the temporary parameters are utilized to compute the testing loss on the meta-testing models. We select $P$ meta-training models and $Q$ meta-testing models, initially calculating the adversarial loss under the meta-training models with the initial encoder parameters $\phi_{en}$:
$\mathcal{L}_{adv}^{tra}=\sum_{p=1}^{P}\mathcal{L}_{adv}(\mathcal{F}_{p}(\hat{x}_{s}),\mathcal{F}_{p}(x_{t})\mid\phi_{en}),$ (8)
where $\mathcal{F}_{p}$ represents the $p$-th meta-training FR model, and $\mathcal{L}_{adv}^{tra}$ represents the obtained meta-training adversarial loss.
@@ 4.2 Muti-Task Compatible Meta-Optimization
4.3 Regularization-Enhanced Information Decoding
The online network processes images uploaded to the platform to conform to the standards for image file storage and transmission. This complex image processing often significantly alters watermarked images, increasing the difficulty of watermark extraction. Therefore, we design a joint noise pool $NP(\cdot)$ that includes differentiable JPEG compression ($QF\!\!=\!\!50$) and Gaussian noise ($Var=0.003$) to enhance the decoder robustness. Watermarked images under both temporary and initial parameters are uniformly processed through the noise pool before decoding. We initially calculate the information loss of the watermarked image obtained under the original encoder parameters using Binary Cross-Entropy (BCE) loss $\mathcal{L}_{BCE}$, as follows:
@@ 4.3 Regularization-Enhanced Information Decoding
The encoder and decoder interact synergistically during co-optimization, serving as mutual priors, as illustrated in Fig. 3. To enhance cross-model gradient compatibility, we introduce an information regularization module. This module calculates information loss using the latest temporary update of the watermarked image, allowing it to further refine the optimization direction of the decoder based on the encoder:
@@
The outer loop optimization of the model is guided by three types of losses: adversarial loss, invisibility loss, and watermark information loss.These losses collectively serve as optimization objectives to simultaneously update the encoder and decoder, searching for the global optimum. The total loss calculation for DIP-Watermark is as follows:
@@
where $\lambda_{adv}$, $\lambda_{en}$ and $\lambda_{wm}$ are positive hyperparameters, and we empirically set them to 100, 0.05 and 0.05, respectively. To balance the significantly different loss values among various tasks and ensure stable training, we set different orders of magnitude for loss weights. During the initial optimization phase, watermark embedding results in severe degradation of the watermark image. At this stage, the invisibility loss measured at the pixel level is considerably larger compared to the adversarial loss. Thus, we address this by assigning smaller weights to the former. Furthermore, we maintain consistency in the orders of magnitude between information loss and image loss weights. This is because, during the initial optimization, the watermark image resembles meaningless noise, and the decoder does not need to learn watermark information from such noise distribution. As the invisibility loss rapidly decreases to the same order of magnitude as the watermark information loss, the watermark image exhibits a learnable pattern distribution, allowing for more balanced and stable updates to the decoder. The impact of weight coefficient settings on performance is demonstrated in the ablation study section.
5 Experiments
5.1 Experimental Settings
5.1.1 Datasets
We evaluate DIP-Watermark on two commonly used large face image datasets: LFW [29] and CelebA [30]. The LFW dataset consists of 13,233 facial images belonging to 5,749 identities. CelebA comprises over 200,000 facial images. We randomly selected 3,000 images from each dataset for watermark model training and 1,000 images for testing. The watermark message is 50 bits, sufficient to meet the requirements for representing identity in the real world. To meet the input size requirements of FR models, all images are initially processed through MTCNN [31] to extract facial regions and sampled to a size of $112\times 112$.
5.1.2 Implementation details
Our DIP-Watermark is implemented by PyTorch and executed on NVIDIA RTX 3090. The entire training process of DIP-Watermark spans 2500 epochs with a batch size of 32. We empirically adjust the Adam optimizer with an initial learning rate of 0.00005 for stable training. We select various offline and online commercial victim FR models to evaluate the black-box transferability of DIP-Watermark,including: 1) Offline models: IR152, IRSE50, MobileFace [26], FaceNet [14], ArcFace [26] based on IResNet100 [32] and CosFace [33] based on IResNet100. 2) Online models: Face++ [34] and Aliyun [35]. To demonstrate the black-box transferability of our method, we alternately selected three models from them as white-box models for integrated training and sequentially obtained black-box test results on the remaining models as in [16, 6, 2], with the parameters set as $P=2$ and $Q=1$.
5.1.3 Baselines
DIP-Watermark is the first watermarking method designed for FR models with both adversariality and traceability, and there is no similar work before. Thus, we first compare its adversariality with adversarial attack methods: 1) Patch-based adversarial attack methods: Adv-Makeup [16], AdvHat [6]; 2) Perturbation-based adversarial attack methods: PGD [17], FGSM [18]. Simultaneously, we straightforwardly combine adversarial adversarial attack methods with watermarking methods Hidden [12] and FakeTagger [20]. We embed watermarks directly into adversarial examples generated by PGD and FGSM, By comparing adversariality and traceability, we demonstrate the advantages of DIP-Watermark, which only requires embedding a single watermark.
  Datasets Type Method IR152 IRSE50 IResNet100-Arc MobileFace FaceNet IResNet100-Cos
@@ 5.1.3 Baselines
LFW Patch-based Adv-Makeup 0.026 N/A 0.236 N/A 0.026 N/A 0.289 N/A 0.026 N/A 0.000 N/A
@@
Watermark-based DIP-Watermark (ours) 0.326 0.992 0.810 0.997 0.381 0.996 0.716 0.997 0.643 0.994 0.355 0.996
@@
Watermark-based DIP-Watermark (ours) 0.433 0.998 0.635 0.998 0.375 0.995 0.679 0.997 0.318 0.995 0.393 0.998
@@
Watermark-based DIP-Watermark (ours) 65.491 0.998 43.945 0.998
@@
5.1.4 Metrics
We use SSIM [36] to evaluate the quality of images. Accuracy $Acc$ is used to evaluate the watermark recovery accuracy of the watermarked image, which is calculated as follows:
$Acc=1-\frac{1}{L}\sum_{l=1}^{L}\left|w_{id}-\hat{w}_{id}\right|,$ (19)
where $L$ denotes the total length of the watermark information and $l$ denotes the index of the watermark bit.
We adopt Attack Success Rate (ASR) [37] for impersonation attack to evaluate adversariality, which is calculated through:
@@ 5.1.4 Metrics
where $i$ represents the image index, $N$ represents the total number of test images, and $\tau$ represents the similarity threshold. The value of $\tau$ will be set as the threshold at 0.01 FAR (False Acceptance Rate) for each victim FR model as most face recognition works do, i.e., IR152 (0.167), IRSE50 (0.241), MobileFace (0.302), FaceNet (0.409), IResNet100-Arc (0.280) and IResNet100-Cos (0.280) [16].
5.2 Comparison with Baselines
To meet the challenges of the practical scenarios, where users are entirely unaware of the underlying structure and the specific training schemes of the FR models used by third-party databases, We benchmark against prior work [16, 8, 1], focusing on evaluating the defense method’s performance against black-box models. For perturbation-based adversarial attack methods, to ensure the generated adversarial images do not affect social usability, we control the average SSIM value of all adversarial images generated by different methods to be 0.9 $\pm$ 0.03 in our experiments for fair comparison. All adversarial attack methods are integrated attacks conducted on same victim models as DIP-Watermark.
As shown in Table II, DIP-Watermark exhibits excellent adversariality and traceability simultaneously. It achieves optimal ASR on almost evaluated black-box FR models, surpassing adversarial example methods that only possess adversariality. Patch-based adversarial attacks generally exhibit poor black-box transferability, particularly with minimal adversarial impact on CosFace and ArcFace models. Notably, Patch-based adversarial attack methods directly alter low-level pixel information, affecting image appearance and social usability. Perturbation-based methods exhibit relatively better black-box transferability; however, all methods consistently demonstrate lower black-box ASR on FaceNet, CosFace, and ArcFace, which can be attributed to the fact that, unlike feature-matching-based FR models, FaceNet with InceptionResnetV1 as the backbone is a classification-based FR model, and the IRSeNet100 shares a similar ResNet variant with it, making transfer attacks more challenging. Nevertheless, DIP-Watermark achieves nearly optimal black-box ASR, especially surpassing perturbation-based methods by up to $50\%$ on FaceNet. While ensuring excellent black-box ASR, DIP-Watermark also maintains superior traceability accuracy. Combined-based methods significantly compromise adversariality after watermark embedding, and embedding watermark information after each image iteration for adversarial example generation drastically reduces computational efficiency. In contrast, DIP-Watermark requires embedding only one robust watermark, with an embedding response time of merely 0.07s, while achieving outstanding black-box ASR and over $99\%$ watermark recovery accuracy.
To assess the practical feasibility of the methods, we employed two commercial online FR APIs for face verification, including Face++ [34] and Aliyun [35]. We upload the adversarial images generated by the different methods with the target facial images to the platforms, and the white-box models used for training are MobileFace, FaceNet, and CosFace. The platforms provided confidence scores indicating whether these two images belong to the same identity. We report the average confidence score (ACS) for 100 images from the LFW dataset. As shown in Table III, DIP-Watermark achieved the highest ACS while maintaining excellent traceability accuracy. This confirms the practical deployability of DIP-Watermark.
  Processing Parameters Type Method IR152 IRSE50 IResNet100-Arc MobileFace FaceNet IResNet100-Cos
@@ 5.2 Comparison with Baselines
Gaussian noise Var=0.003 Patch-based Adv-Makeup 0.026 N/A 0.263 N/A 0.026 N/A 0.289 N/A 0.026 N/A 0.000 N/A
@@
Combination-based PGD$+$Hidden 0.211 0.629 0.424 0.636 0.055 0.632 0.365 0.629 0.033 0.632 0.016 0.632
FGSM$+$Hidden 0.113 0.652 0.145 0.633 0.000 0.636 0.312 0.633 0.011 0.636 0.000 0.641
PGD$+$FakeTagger 0.220 0.982 0.410 0.971 0.023 0.961 0.341 0.971 0.017 0.982 0.022 0.961
FGSM$+$FakeTagger 0.103 0.975 0.312 0.970 0.013 0.963 0.300 0.963 0.011 0.970 0.024 0.963
Watermark-based DIP-Watermark (ours) 0.265 0.975 0.725 0.987 0.308 0.979 0.695 0.987 0.563 0.975 0.311 0.985
JPEG QF=30 Patch-based Adv-Makeup 0.000 N/A 0.052 N/A 0.000 N/A 0.131 N/A 0.000 N/A 0.000 N/A
@@
Combination-based PGD$+$Hidden 0.261 0.667 0.513 0.654 0.032 0.662 0.412 0.654 0.030 0.672 0.045 0.654
FGSM$+$Hidden 0.119 0.686 0.280 0.686 0.000 0.673 0.295 0.692 0.035 0.655 0.016 0.672
PGD$+$FakeTagger 0.260 0.886 0.463 0.892 0.040 0.883 0.440 0.898 0.021 0.883 0.029 0.898
FGSM$+$FakeTagger 0.119 0.887 0.310 0.896 0.016 0.883 0.360 0.887 0.022 0.893 0.011 0.883
Watermark-based DIP-Watermark (ours) 0.272 0.967 0.655 0.958 0.286 0.983 0.736 0.967 0.587 0.983 0.398 0.980
Resize Size=$1/2$ Patch-based Adv-Makeup 0.026 N/A 0.236 N/A 0.026 N/A 0.289 N/A 0.026 N/A 0.000 N/A
@@
Combination-based PGD$+$Hidden 0.241 0.685 0.411 0.691 0.023 0.641 0.444 0.685 0.022 0.663 0.023 0.641
FGSM$+$Hidden 0.109 0.714 0.312 0.762 0.000 0.721 0.333 0.754 0.011 0. 721 0.000 0.736
PGD$+$FakeTagger 0.203 0.987 0.336 0.977 0.025 0.988 0.425 0.987 0.011 0.979 0.016 0.982
FGSM$+$FakeTagger 0.137 0.987 0.246 0.986 0.000 0.982 0.361 0.987 0.011 0.986 0.000 0.986
Watermark-based DIP-Watermark (ours) 0.255 0.990 0.584 0.990 0.211 0.989 0.688 0.990 0.408 0.993 0.253 0.988
Gaussian blur Kernel=3 SD=30 Patch-based Adv-Makeup 0.057 N/A 0.192 N/A 0.057 N/A 0.326 N/A 0.010 N/A 0.002 N/A
AdvHat 0.000 N/A 0.009 N/A 0.000 N/A 0.047 N/A 0.000 N/A 0.000 N/A
Perturbation-based PGD 0.259 N/A 0.440 N/A 0.060 N/A 0.481 N/A 0.042 N/A 0.036 N/A
FGSM 0.196 N/A 0.320 N/A 0.022 N/A 0.361 N/A 0.034 N/A 0.011 N/A
Combination-based PGD$+$Hidden 0.251 0.664 0.403 0.675 0.030 0.664 0.452 0.686 0.033 0.683 0.021 0.672
FGSM$+$Hidden 0.176 0.689 0.306 0.689 0.011 0.672 0.320 0.679 0.010 0. 689 0.000 0.689
PGD$+$FakeTagger 0.272 0.986 0.387 0.983 0.051 0.976 0.416 0.979 0.022 0.968 0.020 0.983
FGSM$+$FakeTagger 0.183 0.981 0.342 0.985 0.000 0.980 0.375 0.973 0.010 0.977 0.023 0.972
Watermark-based DIP-Watermark (ours) 0.296 0.991 0.733 0.987 0.265 0.987 0.668 0.992 0.527 0.988 0.211 0.986
@@
5.3 Robustness Analysis
Online platforms typically process images to conform to standardized transmission and storage protocols. Therefore, the robustness of identity privacy protection schemes needs to be emphasized to ensure their feasibility in practical deployment. We applied Gaussian noise, JPEG compression (JPEG), resizing or Gaussian Blur to the generated images of all methods before evaluating their adversarial robustness and traceability robustness. As shown in Table IV, the ASR of adversarial attack methods significantly decreases after various image processing operations, with a maximum decrease of up to $40\%$ on the IRSE50 model after resizing, while the ASR of DIP-Watermark only decreases by approximately $20\%$, maintaining optimal performance across all evaluated victim models. Furthermore, after image processing, the traceability accuracy of DIP-Watermark consistently remains above 0.96, exceeding that of combined methods by up to $10\%$. This indicates the stable robustness of DIP-Watermark across multiple tasks. This robustness advantage can be attributed to two factors. Firstly, unlike adversarial attack methods based on iterative gradients and direct overlaying of perturbations on images, DIP-Watermark embeds watermark information into the deep feature space of the carrier, with the encoder learning to encode the carrier’s general robust features. Additionally, DIP-Watermark optimizes the watermark encoder and decoder collaboratively while introducing adversarial transferability through meta-optimization strategies, thereby increasing the output dimensions of the model through multitask learning.  [27] has confirmed that multitask learning, especially joint optimization of low-correlation tasks, significantly enhances model robustness.
5.4 Visualization and Analysis
We present the visual results of different methods using samples covering various genders, ages, and degrees of facial occlusion, as shown in Fig. 4. Patch-based methods rely on significant semantic modifications to images. Adv-Makeup has limitations for images with occluded eyes. Perturbation-based methods produce adversarial examples with regular noise patterns, easy to be discovered and defended by adversaries. DIP-Watermark-generated watermarked images exhibit only differences in overall hue, without noticeable artifacts. Additionally, we use Grad-CAM [38] visualization to illustrate the gradient responses of FGSM and DIP-Watermark under black-box and white-box models. As shown in Fig. 5, FGSM and DIP-Watermark both modify gradient responses in the white-box model. While FGSM overfit to modifications in the region of interest of the white-box model, including the background, resulting in poor transferability to black-box models. DIP-Watermark encodes more universal adversarial features of the carrier, focusing on crucial facial regions, and effectively transfers these adversariality to the black-box model.
  Method IR152 IRSE50 IResNet100-Arc MobileFace FaceNet IResNet100-Cos